Length of a rectangle exceeds its width by 5m.If the perimeter of the rectangle is 38 metre,find the length and width of the rectangle?
Answer by Cromlix(4381)

Width = x
Length = x + 5
Perimeter = 2*length + 2*width
38 = 2(x + 5) + 2(x)
38 = 2x + 10 + 2x
38 = 4x + 10
4x = 38 - 10
4x = 28
x = 7m
Width = 7m
Length = 12m
Hope this helps.
:-)